If you deposit money today in an account that pays 12% annual interest, how long will it take to double your money? Round your answer to two decimal places.
Answer #1

Let X be the amount deposited today.

Future value = Present value * (1+ rate per period)^ no. of periods

2X = X*1.12^n

1.12^n = 2

n = log 2 / log 1.12

= 0.30102999566/0.04921802267

= 6.12 years